OpenStudy (help_people):

What is 3 to the power of 2 over 3 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (help_people):

@misty1212

OpenStudy (misty1212):

\[\huge 3^{\frac{2}{3}}\]?

OpenStudy (mathstudent55):

The exponent is 2/3 ?

OpenStudy (misty1212):

\[\huge b^{\frac{m}{n}}=\sqrt[n]{b^m}\]

OpenStudy (anonymous):

HI!

OpenStudy (mathstudent55):

The numerator of the exponent is an exponent, and the denominator of the exponent is the index of a root.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

in a fraction exponent, the numerator is the power, the denominator is the root

OpenStudy (misty1212):

you got that?

OpenStudy (help_people):

yes the exponent is2/3 @mathstudent55

OpenStudy (help_people):

yes i believe so @misty1212

OpenStudy (help_people):

@mathstudent55 may you finsh the question?

OpenStudy (mathstudent55):

When you have a fraction as an exponent, replace the fraction with an exponent and a root. The numerator tells you the exponent. The denominator tells you the root. For example: